### v1

# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

Abstract
Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime. This paper introduces a novel, integrated framework that leverages the power of transformer neural networks and deep reinforcement learning (DRL) algorithms to optimize maintenance actions. Our approach employs the transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the Remaining Useful Life (RUL) of equipment. Simultaneously, the DRL component of our framework provides cost-effective and timely maintenance recommendations. We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions. Consequently, our pioneering approach provides an innovative data-driven methodology for prescriptive maintenance, addressing key challenges in industrial operations and leading the way to more efficient, cost-effective, and reliable systems.
Index Terms:
Prescriptive Maintenance; Transformer; Deep Reinforcement Learning.
I Introduction
The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies. With these advancements, however, comes the heightened need for reliable, efficient, and cost-effective maintenance systems. In particular, as manufacturing processes become increasingly automated, the costs associated with machine downtime rise substantially. Traditionally, human operators have been required to monitor machinery constantly. The COVID-19 pandemic has complicated this scenario, elevating the costs associated with labor-intensive activities and emphasizing the need for reducing on-site staff to minimize infection risks.
Prescriptive maintenance has emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [1]. It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes. According to ISO standard 13381-1 [2], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning.
Current predictive maintenance systems largely rely on heuristic models or conventional machine learning algorithms. While effective to some extent, these approaches often fail to capture complex temporal dependencies in sensor data, limiting their ability to accurately predict RUL or prescribe optimal maintenance actions. Furthermore, many existing solutions require extensive labeled datasets, which are costly and time-consuming to generate.
In response to these challenges, this paper proposes a novel, data-driven framework that leverages the advanced capabilities of Transformer neural networks and DRL algorithms. Transformers excel at modeling long-term dependencies in sequential data and are highly adaptable, making them ideal for handling complex industrial sensor data. DRL algorithms, on the other hand, excel at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks.
The framework aims to achieve the following:
  • Accurately predict the RUL of machinery using Transformer-based models, thus providing early warnings for potential failures.
  • Employ DRL algorithms to offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs.
Unlike existing solutions that address either predictive or prescriptive maintenance, our framework integrates both aspects seamlessly. It leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations. By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges. Additionally, we validate our approach using the NASA C-MPASS dataset, achieving superior performance in both predicting RUL and optimizing maintenance actions, even when compared to traditional neural networks on larger datasets.
Contributions:
  • We are the first to address the unique challenges of limited dataset sizes in the prescriptive maintenance landscape.
  • We propose a DRL-based recommendation system for maintenance actions, reducing both labor involvement and operational costs.
Organization. The rest of the paper is organized as follows: Section II reviews related work in the field of prescriptive maintenance and machine learning algorithms. Section III describes the proposed framework in detail. Section IV presents the experimental setup and results. Finally, Section V concludes the paper and outlines future directions for research.
II Related Work
[…]
Corrective Maintenance is “fault-based maintenance” and is an unplanned mode of maintenance carried out only when equipment malfunctions are detected. Preventive Maintenance, or scheduled inspection and maintenance, generally involves halting the production line according to a predetermined schedule, with an emphasis on repairs before a failure occurs. These two types of maintenance methods are traditionally the most commonly adopted in the industry. However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance. Predictive Maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as Condition-Based Maintenance (CBM). Prescriptive Maintenance is generally considered an extension of Predictive Maintenance, optimizing maintenance strategies based on fault diagnosis results. It emphasizes identifying the inherent reasons for the occurrence of faults and subsequently directing the generation of operational and maintenance strategies [4]. Compared to traditional maintenance methods, Prescriptive Maintenance leverages historical system status to comprehensively forecast future system operations and optimize the resources required for maintenance, aiming to prevent waste from excessive maintenance or production losses due to malfunctions and downtime.
II-A Deep Learning-Based Remaining Useful Life Prediction
Accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations. While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [5, 6]. Notably, they also entail labor-intensive manual feature engineering.
However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data. Deep learning, catalyzed by computational advances, emerges as a remedy. Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [7, 8]. Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
Amongst the deep learning architectures, Transformers, originally conceived for language tasks [9], have recently demonstrated efficacy in time-series data processing. For instance, Zhou et al. [10] integrated a time-embedding layer, augmenting the Transformer’s time-series prediction. Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
II-B Deep Reinforcement Learning enabled Predictive Maintenance
DRL’s foray into predictive maintenance has primarily focused on refining maintenance schedules [11, 1, 12]. While these applications exhibit DRL’s prowess in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, eschewing actionable preventative measures. Gordon et al. [13] did venture into employing DRL for prescriptive maintenance, but the holistic integration of prediction and actionable recommendations remained elusive.
In the annals of maintenance research, preventive and predictive methods have been extensively probed, as highlighted by Ansari [14] and Nemeth [15]. Yet, the burgeoning domain of prescriptive maintenance, transcending mere forecasts to offer optimized interventions, yearns for a unified solution. While Ansari et al. [14] and Nemeth et al. [15] have illuminated this space, the prevalent trend has either skewed towards prediction or prescription. Despite separate advancements in Transformer-based predictions and DRL prescriptive models, a comprehensive framework synergizing these modalities remains absent. This research endeavors to traverse this void, proposing an integrated approach for prescriptive maintenance. By amalgamating the Transformer’s predictive capabilities with DRL’s prescriptive strengths, our framework envisions precise malfunction forecasts paired with optimized action plans.
III Problem Formulation
In the modern industrial ecosystem, machinery and equipment form the backbone of operations. But with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability. The cascade effect is not just limited to production loss; it reverberates through aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [16]. Therefore, the primary challenges reside in not just predicting when machinery might fail but also understanding the myriad of preventive actions that can be employed effectively.
III-A Predicting Remaining Useful Life (RUL)
At the core of predictive maintenance lies the notion of RUL, which provides an estimation of the duration a machine or its components will continue to function before encountering a failure. An accurate RUL prediction empowers industries, enabling them to strategically schedule maintenance activities and, in doing so, avert the unpredicted halting of operations. The RUL offers a dynamic and predictive window into the operational health of a machine. However, deducing this metric isn’t straightforward. The operational data of a machinery component is typically represented as a time series $X=\{x_{1},x_{2},\cdots,x_{t}\}$, where each $x_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions. The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
  • $T$ is the total expected operational lifespan of the machine or component, which might vary based on its usage, environmental factors, and maintenance history.
  • $t$ represents the current age or operational time of the machine.
The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors. The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $X$ to provide an accurate approximation of $R(t)$: $M(X)\approx R(t)$.
III-B Optimal Maintenance Recommendations with DRL
While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions. Upon gaining insights into potential machine downtimes, the challenge shifts to determining the most optimal maintenance action. Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
Consider a machinery component’s state $s$, encapsulating its health, RUL, and other relevant metrics. From this state, we want to choose an action $a$ from a set of potential actions $A$ to maximize the expected future reward. Given a machinery state $s$ which integrates various metrics (like its RUL, health indicators, operational constraints, etc.), the optimal maintenance action $a$ needs to be chosen from a pool of potential actions $A$. The essence of this problem is rooted in the principle of maximizing long-term rewards.
We can mathematically denote a value function $Q(s,a)$, which illustrates the expected cumulative benefit or reward of adopting action $a$ in state $s$. The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in A}Q(s,a)$, where
  • $Q(s,a)$ encompasses immediate benefits and projected future rewards, taking into consideration the dynamic nature of machinery health and operation.
  • The cumulative reward integrates both tangible (like cost savings) and intangible (like averting potential future issues) benefits.
The central challenge here is not only defining the rewards but also ensuring that the decision-making process is robust and can adapt to evolving machinery states and operational requirements.
IV Proposed Framework
Our proposed framework consists of two primary components: a Transformer model for RUL prediction and a subsequent module for recommending maintenance actions. Our proposed framework can be illustrated in the following Algorithm 1 :
Input: Input sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$
Output: Output sequence $\mathbf{y}=(y_{1},y_{2},\ldots,y_{m})$
1 Initialize the positional encoding $\mathbf{PE}$;
2 Initialize the input embeddings $\mathbf{E}$;
3 for $i\leftarrow 1$ to $L$ do
4       Apply multi-head attention with residual connections and layer normalization;
5       Apply position-wise feedforward networks with residual connections and layer normalization;
6
7 end for
8Implement a linear transformation and a softmax activation function;
Input: Observation $s$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
Output: Action $a$
9 Initialize the neural network weights;
10 for each episode do
11       Initialize the starting state $s$;
12       while $s$ is not a terminal state do
13             Select action $a$ using an $\epsilon$-greedy policy;
14             Execute action $a$, observe reward $r$ and successor state $s^{\prime}$;
15             Update neural network weights utilizing the Temporal Difference (TD) error $\delta$ and the learning rate $\alpha$;
16             Adjust the exploration rate $\epsilon$;
17             Update current state $s\leftarrow s^{\prime}$;
18
19       end while
20
21 end for
Algorithm 1 The Transformer and Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework
IV-A Remaining Useful Life Prediction Module
Transformers, rooted in Natural Language Processing with models like BERT, GPT-2, and T5 [17], have become valuable tools in time-series data analysis. This is largely due to their proficiency in understanding complex temporal sequences. Our model harnesses the Transformer’s capabilities to forecast the Remaining Useful Life (RUL) from historical time-series sensor data. At the heart of this approach lies the self-attention mechanism, which prioritizes different time steps based on their importance, adapting finely to the intricacies of the data.
IV-A1 Encoder Module
For a given time-series $x=\{x_{1},x_{2},...,x_{T}\}$, the Encoder Module converts it into hidden states $h=\{h_{1},h_{2},...,h_{T}\}$. Each Transformer block in this module comprises:
  • Multi-Head Self-Attention Mechanism: Given matrices for query $Q$, key $K$, and value $V$, the attention mechanism is:
    $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$
    with $d_{k}$ being the dimension of queries and keys.
  • Feed-Forward Neural Network: To capture non-linear patterns, this is represented as:
    $FFN(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2},$ (1)
    with weight matrices $W_{1},W_{2}$ and bias terms $b_{1},b_{2}$.
IV-A2 Self-Attention Mechanism
This module facilitates the model in emphasizing or downplaying certain temporal characteristics. The attention weights $\alpha$ compute the weighted sum of features:
$\alpha_{ij}=\text{Softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}\right),$
where $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
IV-A3 Decoder Module
Post the encoding of time-series data into hidden states, the decoder refines this information to produce outputs conducive for RUL prediction. Given hidden states $h$, the Decoder Module yields output states $o=\{o_{1},o_{2},...,o_{T}\}$. Each block in the decoder encompasses:
  • Multi-Head Cross-Attention Mechanism: Attending to the encoder’s outputs while forming the decoder’s states, it’s represented as:
    $\text{Attention}(Q_{d},K_{e},V_{e})=\text{Softmax}\left(\frac{Q_{d}K_{e}^{T}}{\sqrt{d_{k}}}\right)V_{e},$
    with $d_{k}$ being the dimension of the queries and keys.
  • Feed-Forward Neural Network: To discern patterns in the decoder states:
    $FFN(o)=\max(0,oW_{3}+b_{3})W_{4}+b_{4},$
    with weight matrices $W_{3},W_{4}$ and biases $b_{3},b_{4}$.
IV-A4 Prediction Module
At the culmination of our model, the prediction module utilizes the hidden states $h$ to estimate the RUL. This involves one or several dense layers, expressed as:
$\text{RUL}=\sigma(hW+b),$
with weight matrix $W$, bias $b$, and activation function $\sigma$ ensuring the RUL remains within the intended range.
IV-B Maintenance Action Recommendation Module
Upon retrieving the RUL predictions from the previous module, the maintenance action recommendation module initiates. This module uses the predicted RUL as a primary determinant, facilitating the calibration of the urgency, type, and extent of necessary maintenance interventions. However, an extended predicted RUL diminishes the immediate demand for aggressive maintenance interventions. In contrast, a reduced RUL necessitates prompt preventive or remedial actions. Recommendations from this module undergo further refinement, adapting to specific industrial settings and the complexities inherent to the particular equipment or system.
This module predominantly employs two methodologies:
  • Rule-Based Systems: These systems are grounded in predetermined heuristic principles. As an instance, should the RUL prediction for a component fall below a designated limit (e.g., 10 hours), an immediate maintenance notification is dispatched. On the other hand, if the RUL surpasses a set benchmark, periodic inspections may be deemed adequate.
  • Machine Learning Classifiers: These classifiers are cultivated using historical datasets, where earlier RUL predictions are coupled with their corresponding effective maintenance interventions. Inputting the present RUL prediction enables the classifier to recommend the most probable efficacious maintenance strategy, grounded in historical insights.
The harmonious integration of the RUL predictions with the Maintenance Action Recommendation Module ensures that maintenance schedules are both anticipatory and informed by data, culminating in augmented system dependability, extended equipment longevity, and streamlined operational expenditures.
Upon acquiring the RUL predictions from the antecedent module, the Maintenance Action Recommendation Module becomes operational. The forecasted RUL, indicative of a system’s operational state, is assimilated into the state representation within the Markov Decision Process (MDP) framework. Subsequently, this representation guides the meticulous formulation of optimal maintenance strategies over time.
IV-B1 Markov Decision Process (MDP) for Maintenance
  • Environment State Space. As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S_{t}}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ encapsulating all possible states the system can take at time $t$.
  • Action Space. If we denote the number of possible maintenance actions by $m$, then the action space $A$ can be represented as $A=\{a_{1},a_{2},\ldots,a_{m}\}$.
  • State Transition Probabilities. The probability of transitioning from state $s_{i}^{t}$ to state $s_{j}^{t+1}$ given action $a_{k}^{t}$ is represented by:
    $P_{i,j}^{k}=P(s_{t+1}=s_{j}^{t+1}|s_{t}=s_{i}^{t},a_{t}=a_{k}^{t}).$
    The complete state transition probability matrix given action $a_{k}^{t}$ would then be a matrix $\mathbf{P^{k}}$ where each entry $P_{i,j}^{k}$ represents the transition probability from state $s_{i}^{t}$ to state $s_{j}^{t+1}$.
  • Reward Structure. The immediate reward received after transitioning from state $s_{i}^{t}$ to $s_{j}^{t+1}$ upon taking action $a_{k}^{t}$ is given by $R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1})$. The expected reward for taking action $a_{k}^{t}$ in state $s_{i}^{t}$ is:
    $R(s_{i}^{t},a_{k}^{t})=\sum_{j=1}^{n}P_{i,j}^{k}R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1})$
  • Policy. A policy $\pi$ is a mapping from states to actions. An optimal policy $\pi^{*}$ maximizes the expected cumulative discounted reward from any starting state. The value of a state $s_{i}$ under policy $\pi$ is:
    $V^{\pi}(s_{i})=R(s_{i},\pi(s_{i}))+\gamma\sum_{j=1}^{n}P_{i,j}^{\pi(s_{i})}V^{\pi}(s_{j}),$
    where $\gamma$ is the discount factor.
  • Bellman Optimality Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the recursive relationship:
    $V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$
V Experiments
In this section, we evaluate our proposed framework on the real-world dataset. Experiments are done on the XXX.
V-A Dataset
The NASA’s C-MAPSS dataset [18] is partitioned into four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes. The training data comprises inputs from three operating conditions and 21 sensors’ measurements. For training, every engine data point serves as a sample paired with its RUL. We applied a piecewise linear degradation model [19] for RUL determinations. In testing, the last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
V-B Evaluation Metric
To assess the model’s predictive accuracy, we utilized the Mean Squared Error (MSE), given by
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$ (2)
which quantifies the average squared differences between predicted $\hat{y}_{i}$ and observed $y_{i}$ values. Additionally, the coefficient of determination, $R^{2}$, defined as
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (3)
where $\bar{y}$ is the mean of observed values, was employed to measure the proportion of variance explained by the model. A higher $R^{2}$ suggests a better fit to the observed data.
V-C Remaining Useful Life Prediction
The Temporal Fusion Transformer (TFT) [20] addresses the limitations of traditional time series forecasting models by offering enhanced interpretability and the capability to deal with multiple prediction horizons. It merges the advantages of Recurrent Neural Networks (RNNs) and Transformers [21] into a unified encoder-decoder architecture.
Encoder: The encoder transforms historical time-series data into a high-dimensional space using a blend of convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
Temporal Attention: This specialized module empowers the model to focus on different time-steps variably, facilitating the capture of dynamic temporal patterns and long-range dependencies.
Decoder: Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for subsequent steps.
Static Inputs: Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
Interpretable Components: To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
Figure 1: FD001.
Figure 2: FD002.
Figure 3: FD003.
Figure 4: FD004.
V-D Optimal Maintenance Actions Recommendation
In the previous stage, we succeed in predicting RUL using four datasets from NASA C-MAPSS. But results from the prediction cannot provide enough information for engineers to make decisions on which maintenance actions to take. Thus, we may recommend corresponding actions based on the cost and benefits. To optimize results, we leverage DRL algorithms, which helps to find the best trade-off.
V-D1 States and Actions
The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset. These PCA features provide a compact yet informative snapshot of the machine’s current health condition. The action space contains three primary actions:
  • No Action ($a_{1}$), appropriate for healthy machine states, minimizes immediate costs but risks penalties in critical conditions;
  • Partial Maintenance ($a_{2}$), suitable for moderate wear states, involves preventive actions like lubrication to extend machine life at moderate costs;
  • Complete Overhaul ($a_{3}$), necessary for critical machine states, entails comprehensive repairs and replacements to avert catastrophic failures but incurs high costs and downtime.
The DRL agent aims to navigate this state-action space to minimize long-term operational costs by selecting the optimal maintenance actions based on each state’s representation of machine health.
V-D2 Probabilities Model
In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset. Specifically, each entry in the state transition matrix $p(S_{t+1}|S_{t},a_{t})$ reflects the probability of transitioning from state $S_{t}$ to $S_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset. Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset. In this MDP framework, the transition matrix has dimensions $(10,3,10)$, representing 10 states, 3 actions, and 10 possible next states, all closely mirroring scenarios that commercial aero-engines might encounter.
V-D3 Reward Model
In predictive maintenance applications using DRL, the objective is not only to maximize the RUL of machinery but also to minimize the operational costs and downtime. To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner. The reward function $R(S_{t},a_{t},S_{t+1})$ is formulated as:
$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})~{}$
$-\beta\times\text{Cost}(a_{t})-\gamma\times\text{DownTime}(a_{t}),$ (4)
where
  • $\text{RUL}_{\text{gain}}(S_{t},a_{t})$ represents the gain in RUL as a consequence of taking action $a_{t}$ at state $S_{t}$. This gain is computed using the state-transition probabilities $p(S_{t+1}|S_{t},a_{t})$ derived from the C-MAPSS dataset. A higher RUL gain is desirable as it suggests an extension in the machine’s operational lifespan.
  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on real-world operation costs and directly affects the overall expenditure.
  • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$. Downtime is critical in operational settings where machine availability significantly impacts productivity or service levels.
  • $\alpha,\beta,\gamma$ are hyperparameters that weigh the importance of RUL, cost, and downtime, respectively. By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL of machinery, minimizing costs, or reducing downtime, as per the organizational objectives.
V-D4 Hyperparamters
To train the algorithm, we use several hyperparameters that were carefully chosen based on empirical experiments. In the following, we present detailed hyperparameters for each DRL algorithm.
Deep Q-network (DQN). DQN alleviates the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function. It also introduces experience replay to improve learning stability and efficiency. The learning rate was set to $0.0001$, the discount factor was set to $0.99$, and the replay buffer size was set to $1,000,000$ experiences. We used a batch size of $512$ and updated the target network every $50$ steps. The exploration vs. exploitation trade-off was controlled by an epsilon value that started at $1.0$ and decayed over time to a final value of $0.01$. Additionally, we used a neural network with two hidden layers of $256$ units each, and we used the Adam optimizer with a learning rate of $0.0001$.
Proximal Policy Optimization (PPO). PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy. We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$. A clip range of $0.2$ was used to limit the updates to the policy, and we applied a value function coefficient of $0.5$ to balance the importance of the policy and the value function. Additionally, we set the entropy coefficient to $0.01$ to encourage exploration, and we used a batch size of $64$. We optimize the policy every 10 epochs and used the Adam optimizer with an epsilon value of $0.00001$. Finally, we set the GAE parameter to $0.95$ to estimate the advantage function.
Soft Actor-Critic (SAC). SAC is particularly effective for problems with high-dimensional action spaces. SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient. The learning rate for the actor and critic networks was set to $0.0003$, and we used a discount factor of $0.99$. We applied a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function. The target entropy was set to $-2.0$, and we used a replay buffer of $1,000,000$ experiences with a batch size of $256$. We also applied a soft update of the target networks with a rate of $0.005$, and we initialized the log standard deviation of the policy with a value of $-2$. We trained the algorithm for $100,000$ time steps and used the Adam optimizer with an epsilon value of $0.0001$.
V-E Results and Discussion
We compare the performance of DQN [22], PPO [23], and SAC [24] on our designed task in the Figure 5, Figure 6, and Figure 7. The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes. The experimental results show that PPO outperforms both DQN and SAC, achieving higher total rewards on the task. The average total rewards obtained by PPO are $55,000$, while DQN and SAC achieve average total rewards of $50,000$ and $45,000$, respectively. However, DQN converges faster than PPO and SAC.
In general, PPO and SAC are considered to be more sample-efficient than DQN, which means that they can achieve comparable performance with fewer samples. This is because PPO and SAC use more efficient updates that make better use of the available data. In contrast, DQN can be more sensitive to the choice of hyperparameters and the amount of data used during training, which can affect its convergence speed.
However, the performance of these algorithms can vary depending on the specific task and the quality of the implementation. In some cases, DQN may converge faster than PPO or SAC, especially if the action space is discrete and the observations are low-dimensional. On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster. Results that are shown in Figure 5 and Figure 7 have confirmed this theory.
Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve the optimal performance.
Figure 5: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
Figure 6: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
Figure 7: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in Cartpole environment.
VI Conclusions
In this paper, we have presented a comprehensive prescriptive maintenance framework that uniquely integrates real-time RUL estimations with reinforcement learning algorithms for dynamic action recommendations. The framework starts by employing the transformer algorithm to estimate the RUL distribution of engines, offering a robust, adaptive approach to understanding machine health in real-time. As machinery operates, our model continues to dynamically update these RUL estimates, providing an ever-evolving understanding of the equipment’s condition.
We rigorously compared the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates. Our findings indicate that the transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process.
For recommending maintenance actions, we used SAC, DQN, and PPO. Each of these algorithms has its strengths and weaknesses, but they all were adapted to work seamlessly with the dynamically updated RUL estimates. Our experimental results showed that SAC excelled in scenarios where the cost function was highly non-linear, while PPO and DQN were more suited for environments with more predictable dynamics.
This framework opens up new avenues for research, particularly in the fine-tuning of hyperparameters that govern the trade-offs between extending the machinery’s useful life and minimizing operational costs. Future work could involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machinery.
In conclusion, our work signifies a step forward in the quest for more intelligent, adaptive, and cost-effective prescriptive maintenance systems. By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a holistic solution that dynamically adapts to the operational realities, thereby driving both economic and operational efficiencies.