### v2

# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

Abstract
Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime. This paper introduces an integrated framework that leverages the capabilities of the Transformer model-based neural networks and deep reinforcement learning (DRL) algorithms to optimize system maintenance actions. Our approach employs the Transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the remaining useful life (RUL) of an equipment. Additionally, the DRL component of our framework provides cost-effective and timely maintenance recommendations. We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions, compared to the other prevalent machine learning-based methods. Our proposed approach provides an innovative data-driven framework for industry machine systems, accurately forecasting equipment lifespans and optimizing maintenance schedules, thereby reducing downtime and cutting costs.
Index Terms:
Prescriptive Maintenance; Transformer; Deep Reinforcement Learning.
I Introduction
The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies [1]. With these advancements, however, comes the increased requests for reliable, efficient, and cost-effective maintenance systems. Prescriptive maintenance has thus emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [2]. It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes. According to ISO standard 13381-1 [3], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning. However, current predictive maintenance systems largely rely on heuristic models for action planning or conventional machine learning algorithms for sensor data processing. While effective to some extent, existing approaches often fail to capture complex temporal dependencies in sensor data due to issues like vanishing or exploding gradients, limiting their ability to accurately predict Remaining Useful Life (RUL).
Apart from that, there is a trend in recent works to leverage Reinforcement Learning (RL) to make plans in maintenance tasks [4]. RL excels at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks. As a result, the RL-based decisions offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs. However, these methods consistently encounter the cold start problem [5], coupled with low learning efficiency. In some special cases such as aeroengine, these methods also suffer from a lack of reliability and explainability. This often leads to their decisions being perceived as untrustworthy by users.
Unlike existing solutions that address either predictive or prescriptive maintenance, we propose a framework named TranDRL, which integrates both aspects seamlessly. TranDRL leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations. By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges by Transformer neural networks and Reinforcement Learning with Human Feedback (RLHF) algorithms. Transformers excel at modeling long-term dependencies in sequential data by the attention mechanism and are highly adaptable, making them ideal for handling complex industrial sensor data. By leveraging the expertise and insights of human operators or maintenance professionals, the algorithm can efficiently guide its actions, accelerating the convergence of the DRL algorithm towards optimal maintenance strategies. This human-guided approach ensures that the algorithm remains on the right track, continually improving its performance and ultimately leading to more efficient and cost-effective industrial maintenance practices.
Contributions. The contributions of this paper can be summarized as follows:
  • To improve the accuracy of the RUL prediction module, we adopt Transformer to address the difficulty in reflecting long-term dependencies of sequential sensor data in industrial scenarios.
  • We propose a DRL-based recommendation system for determining maintenance actions, reducing both labor involvement and operational costs. Furthermore, we utilize RLHF to confirm the correct maintenance actions and improve learning efficiency.
  • A Transformer-DRL framework is proposed to optimize the prescriptive maintenance process, which can be extended to multiple manufacturing scenarios as well.
  • Extensive simulation experiment results, conducted on different datasets, demonstrate that our method outperforms baseline methods in terms of both predicting RUL and optimizing maintenance actions.
Organizations. The rest of the paper is organized as follows. Section II reviews related work in the field of prescriptive maintenance utilizing machine learning algorithms. Section III formulate the problem we are trying to solve. Then, section IV describes our proposed framework for Transformer-driven prediction maintenance in detail. Next, section V presents a case study on our proposed framework using a real-world dataset. Finally, section VI concludes the paper and outlines future directions for research. Notations used across the paper is provided in Table I.
TABLE I: Notations.
Symbol Description
$\mathbf{X}$
Time series of sensor readings and operational parameters.
$R(t)$
Remaining Useful Life at time $t$: $R(t)=T-t$.
$T$
Total expected lifespan of the component.
$t$
Current operational time of the machine.
$M$
Model estimating $R(t)$ from $\mathbf{X}$.
$s$
State encapsulating health and RUL metrics.
$\mathcal{A}$
Set of maintenance actions.
$a$
Action chosen to maximize future reward.
$a^{*}$
Optimal action maximizing cumulative benefit.
$Q(s,a)$
Expected reward of action $a$ in state $s$.
$\epsilon$
Exploration rate in DRL.
$\alpha$
Learning rate in DRL.
$\gamma$
Discount factor for future rewards.
$\bm{y}$
Predicted RUL output by Transformer model.
$\bm{s}$
Observation state in DRL from RUL predictions.
$d_{\text{model}}$
Input/output dimension in Transformer.
$d_{k},d_{v}$
Dimensions of keys/queries and values in attention.
$\mathbf{Q},\mathbf{K},\mathbf{V}$
Query, key, and value matrices in attention.
$FFN(\cdot)$
Feed-forward network in Transformer.
$\bm{\Lambda}$
Attention weight matrix.
$\alpha_{ij}$
Attention weight from $j$-th input to $i$-th output.
$\bm{h}$
Hidden states in encoder.
O
Output states from decoder.
$\sigma(\cdot)$
Activation function in prediction module.
$K$
Number of machines in federated learning setup.
$f_{\text{ensemble}}$
Function for RUL prediction in multi-machine case.
$\mathbf{W},b$
Weight matrix and bias in prediction module.
PCA
Used for selecting important parts of RUL.
Figure 1: An Overview of TranDRL Framework.
II Related Work
[…]
Corrective maintenance is “fault-based maintenance” and is an unplanned mode of maintenance carried out only when equipment malfunctions are detected. Preventive maintenance, or scheduled inspection and maintenance, generally involves halting the production line according to a predetermined schedule, with an emphasis on repairs before a failure occurs. These two types of maintenance methods are traditionally the most commonly adopted in the industry. However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance. Predictive maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as condition-based maintenance. Prescriptive maintenance is generally considered an extension of predictive maintenance, optimizing maintenance strategies based on fault diagnosis results. It emphasizes identifying the inherent reasons for the occurrence of faults and subsequently directing the generation of operational and maintenance strategies [7]. Compared to traditional maintenance methods, prescriptive maintenance leverages historical system status to comprehensively forecast future system operations and optimize the resources required for maintenance, aiming to prevent waste from excessive maintenance or production losses due to malfunctions and downtime.
II-A Deep Learning-Based Remaining Useful Life Prediction
In production maintenance tasks, accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations. While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [8, 9]. Notably, they also entail labor-intensive manual feature engineering.
However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data. Deep learning, boosted by computational advances, emerges as a remedy. Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [10, 11]. Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
Amongst the deep learning architectures, Transformers, originally conceived for language processing tasks [12], have recently demonstrated efficacy in time-series data processing. For instance, Zhou et al. [13] integrated a time-embedding layer, augmenting the Transformer’s time-series prediction. Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
II-B Deep Reinforcement Learning enabled Predictive Maintenance
DRL’s concept into predictive maintenance has primarily focused on refining maintenance schedules [14, 2, 4]. While these applications exhibit DRL’s capability in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, focusing on actionable preventative measures. Gordon et al. [15] explore into employing DRL for prescriptive maintenance, but the integration of prediction and actionable recommendations remained elusive.
In existing maintenance studies, we have explored a lot at preventive and predictive methods, as mentioned in [16] and [17]. But now, there is growing interest in prescriptive maintenance, which goes beyond just predicting problems to suggesting the best actions to take. Even though the work in [16] and [17] has started to light the way, most efforts are either focused on predicting what will happen or on prescribing what to do. Even with improvements in using Transformer technology for predictions and DRL for making recommendations, there is no plan to integrate these two ideas together. Our study aims to fill this gap by introducing a way to use both the forecasting power of Transformers and the decision-making abilities of DRL to not only predict equipment failures but also suggest the best actions to prevent them.
II-C Deep Reinforcement Learning with Human Feedback
Inspired by preference theory [18], Reinforcement Learning from Human Feedback (RLHF) improves the performance of DRL by leveraging feedbacks from human, and releases the difficulty of designing reward function by incorporating human guidance and expertise [19]. To tackle the problem of informative limitation and expensive time consumption, Ibarz et al.  [20] combine expert demonstrations and trajectory preferences together and attempt to let human interact with agent directly. To enhance sample efficiency, other follow-up works leverage semi-supervised learning and data augmentation techniques into RLHF [21, 22]. Most previous works are under the non-Markov assumption which poses a threat to modeling reward function, Early et al.  [23] release this assumption and formulate the scenario into a multiple instance learning process. With the prosperity of large language model recently, RLHF also shows its power in fine-tuning step before the model development in this specific domain [24, 25]. Furthermore, in safety-critical applications like aeroengine maintenance, deviating from standard procedures can have grave consequences. Such non-standard maintenance practices might result in serious safety incidents. Combining human feedback can also help in setting safe boundaries and ensuring that the agent’s behavior remains within acceptable limits. To make the policy more reliable, the incorporation of human feedback can make the agent’s behavior more predictable and better aligned with human expectations, enhancing the explainability of its actions and building trust among users.
III Problem Formulation
In the modern industrial ecosystem, machinery and equipment form the backbone of operations. However, with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability. The cascade effect is not just limited to production loss; it also affects aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [26]. Therefore, the primary challenge is to recommend proper maintenance actions effectively based on accurate prediction of failures of the machine.
III-A Predicting Remaining Useful Life (RUL)
Predicting RUL is essential for predictive maintenance, which provides an estimation of the duration that a machine or its components will continue to function before encountering a failure. An accurate RUL prediction enable the industries to strategically schedule their maintenance activities. The RUL offers a dynamic and predictive window into the operational health of a machine. The operational data of a machinery component is typically represented as a time series $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, where each $\bm{x}_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions. The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
  • $T$ is the total expected operational lifespan of the machine or component, which can vary based on its usage, environmental factors, and maintenance history.
  • $t$ represents the current age or operational time of the machine.
The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors, for example, machine conditions. The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $\mathbf{X}$ to provide an accurate approximation of $R(t)$: $M(\mathbf{X})\approx R(t)$.
III-B Optimal Maintenance Action Recommendations with DRL
While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions. Upon gaining prediction results of potential machine downtimes, the challenge shifts to determining the most optimal maintenance action. Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
Consider a machinery component’s state $s$, which encapsulates its health, RUL, and other relevant metrics. From the state, we want to choose an action $a$ from a set of potential actions set $\mathcal{A}$ to maximize the expected future reward.
The essence of this problem is rooted in the principle of maximizing long-term rewards. We denote a value function $Q(s,a)$, which characterizes the expected cumulative benefit or reward of adopting action $a$ in state $s$. The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, where
  • $Q(s,a)$ encompasses immediate benefits and projected future rewards, taking into consideration the dynamic nature of machinery health and operation.
  • The cumulative reward should integrate both current and weighted future’s benefits strategically.
IV Proposed Framework
To solve the problem above, we propose a framework consisting of two major components: a transformer-based learning model to predict the RUL and a decision module to recommend maintenance actions as illustrated in Figure 1 and Algorithm 1. Specifically, we consider two cases for the RUL prediction, including a single machine case and a multi-machine case.
Input: Sequence of sensor readings $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
Output: Recommended maintenance action $a$
/* Phase 1: RUL prediction using Transformer framework */
1 Initialize the Transformer model with appropriate hyperparameters;
2 Process input sequence $\mathbf{X}$ through Transformer layers;
3 Generate RUL prediction $\bm{y}=(y_{1},y_{2},\ldots,y_{m})$;
/* Phase 2: Maintenance action recommendation using DRL */
4 Initialize DRL model with neural network weights;
5 Use RUL prediction $\bm{y}$ as the observation state $\bm{s}$;
6 Select action $a$ based on $\epsilon$-greedy policy using state $\bm{s}$;
7 Execute action $a$ and observe new state and reward;
8 Update DRL model using observed reward and new state;
9 Adjust $\epsilon$ for exploration-exploitation balance;
10 Output recommended action $a$;
return $a$
Algorithm 1 A Transformer-enabled DRL framework for maintenance selection.
IV-A Case 1: Remaining Useful Life Prediction Module for a Single Machine
Transformers, rooted in natural language processing [27], have become valuable tools in time-series data analysis. This is largely due to their proficiency in understanding complex temporal sequences. Our model leverages the Transformer’s capabilities to forecast the RUL from historical time-series sensor data. The core of this approach is the self-attention mechanism, which establishes causal relations between different time steps based on query-and-key similarities, adapting to the complexity of the data, for example, self-attention can handle complex relationships in machinery sensor data when issues depend on small changes over time in predictive maintenance.
IV-A1 Encoder Module
For a given time-series $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]$, where $n$ is the sequence length and $\bm{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$, the Encoder Module converts $\mathbf{X}$ into hidden states $\bm{h}=[h_{1},h_{2},\cdots,h_{n}]$, with each $h_{i}\in\mathbb{R}^{d_{\text{model}}}$. Each Transformer block in this module comprises:
  • Multi-Head Self-Attention Mechanism. Given matrices for query $\mathbf{Q}\in\mathbb{R}^{n\times d_{k}}$, key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$, and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$, the attention mechanism is:
    $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$
    with $d_{k}$ being the dimension of queries and keys, and $d_{v}$ the dimension of values.
  • Feed-Forward Neural Network. To capture non-linear patterns, this is represented as:
    $FFN(\bm{x}_{i})=\max(0,\bm{x}_{i}\mathbf{W}_{1}+b_{1})\mathbf{W}_{2}+b_{2},$ (1)
    with weight matrices $\mathbf{W}_{1},\mathbf{W}_{2}$ and bias terms $b_{1},b_{2}$.
IV-A2 Self-Attention Mechanism
This module facilitates the model in emphasizing or downplaying certain temporal characteristics. The attention weight $\bm{\Lambda}=[\alpha]_{i,j}$ computes the weighted sum with its $(i,j)$-th element defined as:
$\alpha_{ij}=\text{Softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{j}^{T}}{\sqrt{d_{k}}}\right),$ (2)
where $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
IV-A3 Decoder Module
After encoding the time-series data into hidden states, the decoder further refines this information to produce outputs which are conductive for RUL prediction. Given hidden states $\bm{h}^{\prime}$, the decoder module generates output states $\textbf{O}=\{\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}\}$. Each block in the decoder encompasses:
  • Multi-Head Cross-Attention Mechanism. Cross-attention takes in the encoder’s outputs and combines them with the decoder’s states. This process is represented as
    $\text{Attention}(\mathbf{Q}_{d},\mathbf{K}_{e},\mathbf{V}_{e})=\text{Softmax}\left(\frac{\mathbf{Q}_{d}\mathbf{K}_{e}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}_{e},$ (3)
    where $d_{k}$ represents the dimension of the queries and keys.
  • Feed-Forward Neural Network. This module extracts patterns in the decoder states:
    $FFN(\bm{o}_{i})=\max(0,\bm{o}_{i}\mathbf{W}_{3}+b_{3})\mathbf{W}_{4}+b_{4},$
    with weight matrices $\mathbf{W}_{3},\mathbf{W}_{4}$ and biases $b_{3},b_{4}$.
IV-A4 Prediction Module
The prediction module utilizes the states $\bm{h}$ to estimate the RUL. This involves one or several dense layers, which can be expressed as
$\text{RUL}=\sigma(\bm{h}\mathbf{W}+b),$
where the weight matrix $\mathbf{W}$, bias $b$, and activation function $\sigma$ ensure that the RUL remains within the range.
Input: Input sequence $\mathbf{X}=(\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{n})$
Output: Output sequence $\mathbf{y}=(y_{1},y_{2},\ldots,y_{m})$
1 Initialize the positional encoding;
2 Initialize the input embeddings;
3 for $i\leftarrow 1$ to $L$ do
       /* Encoder Layer $i$/
4       Apply multi-head self-attention to the input sequence;
5       Add residual connection followed by layer normalization;
6       Apply position-wise feed-forward networks;
7       Add another residual connection followed by layer normalization;
8
9 end for
/* Decoder */
10 if task requires a decoder then
11       for $j\leftarrow 1$ to $L$ do
             /* Decoder Layer $j$/
12             Apply masked multi-head self-attention to the output of the encoder;
13             Add residual connection followed by layer normalization;
14             Apply multi-head attention combining encoder outputs and decoder inputs;
15             Add residual connection followed by layer normalization;
16             Apply position-wise feed-forward networks;
17             Add another residual connection followed by layer normalization;
18
19       end for
20
21 end if
22Generate output sequence $\mathbf{y}$ based on the final layer’s output;
return $\mathbf{y}$
Algorithm 2 A Transformer algorithm for a single machine.
IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
The multi-machine case is featured by a federated model, which dispatches an individual training model for similar machines. The local model takes as input the local time series data $\mathbf{X}$ and outputs an estimation of the RUL $R(t)$. By combining the strengths of multiple models extracted from different machines, the approach improves the reliability and accuracy of RUL predictions, enabling industries to make informed decisions regarding maintenance schedules and machinery operations. The averaged model for RUL prediction leverages an ensemble of predictive models to provide a more precise estimate of RUL, enhancing the efficiency and reliability of predictive maintenance in industries.
As shown in the Algorithm 3, given a set of time series data $\mathbf{X}$, the ensemble model aims to predict the RUL $R(t)$ of multiple machines. Formally, the ensemble model can be defined as a function:
$R(t)=f_{\text{ensemble}}(\mathbf{X}),$ (4)
where $f_{\text{ensemble}}$ is the ensemble model that aggregates the predictions of multiple individual models.
By integrating the predictions of multiple models, the ensemble approach refines the reliability and accuracy of the RUL predictions. This enhanced prediction quality allows industries to make precise maintenance schedules and optimize machine operations. The averaged model for RUL prediction can be represented as:
$R(t)=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{X}),$ (5)
where $m$ is the number of models in the ensemble, and $f_{i}$ represents the individual predictive model. This averaging method capitalizes on the collective strength of diverse predictive algorithms. Based on the predictive RULs, the infrastructure can conduct equipment maintenance or replacement to ensure uninterrupted and reliable functionality of the machines.
Input: Distributed input sequences $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{K}$ from $K$ machines
Output: Aggregated output sequences $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
1 Initialize a central model with positional encoding and input embeddings; for each training round do
       /* Local Training on Each Machine */
2       for each machine $k=1$ to $K$ do
3             Receive model parameters from the central model;
4             Apply Algorithm 2 with the given number of layers on local sequence $\mathbf{x}_{k}$;
5             Compute local updates and send them back to the central model;
6
7       end for
      /* Aggregation at the Central Model */
8       Aggregate local updates from all machines;
9       Update central model parameters;
10
11 end for
/* Output Generation */
12 for each machine $k=1$ to $K$ do
13       Send final model parameters to machine $k$;
14       Generate output sequence $\mathbf{y}_{k}$ using the final model on local input $\mathbf{x}_{k}$;
15
16 end for
return $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
Algorithm 3 A Federated Transformer Algorithm for Multiple Machines
The distinction between Case 1 and Case 2 lies primarily in their scale and data handling methods. Case 1 focuses on predicting the RUL for a single machine using a Transformer model, which excels in interpreting complex temporal sequences in sensor data. This approach is highly specialized for analyzing the intricate patterns of an individual machine. In contrast, Case 2 extends to multiple machines, employing a federated learning approach with an ensemble model. This method aggregates data from various machines, synthesizing diverse inputs for a broader analysis. It suits for a industry which contains multiple similar production lines or machines. Additionally, Case 2 could apply Case 1’s algorithm to each machine and then focus on multi-machine data aggregation.
IV-C Maintenance Action Recommendation Module
Upon obtaining the RUL predictions from the previous module, the maintenance action recommendation module initiates. This module uses the predicted RUL as a primary determinant, facilitating the calibration of the urgency, type, and extent of necessary maintenance interventions. If the predicted RUL of a machine or component is longer, there is less need for urgent maintenance. In contrast, if the predicted RUL is shorter, it is important to quickly perform maintenance actions or repairs to prevent problems. Thus, we select the most important part of RUL using principal component analysis (PCA) as follows:
Rule-Based Systems. These systems are established on the basis of predetermined heuristic principles. For instance, should the RUL prediction for a component fall below a designed limit (e.g., 10 hours), an immediate maintenance notification is dispatched. On the other hand, if the RUL surpasses the limit, periodic inspections is considered sufficient.
Strategy-Learning Agents. These agents are trained using historical datasets, where earlier RUL predictions are coupled with their corresponding effective maintenance interventions. Inputting the present RUL prediction enables the agent to recommend the most probable efficient maintenance actions, grounded in historical insights.
The appropriate integration of the RUL predictions with the maintenance action recommendation module ensures that maintenance schedules is efficient and data-driven, ultimately leading to improved system reliability, longer equipment lifespan, and reduced operational costs. Thus, the predicted RUL, indicative of a system’s operational state, is assimilated into the state representation within a Markov Decision Process (MDP)-based framework. The subsequent discussions elaborate the maintenance strategy-learning procedure under the MDP framework, which includes following main components:
Environment State Space. As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S}^{t}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ consisting all possible $n$ states that the system can take at time $t$.
Action Space. If we denote the number of possible maintenance actions by $m$, the action space $\mathcal{A}$ can be represented as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{m}\}$.
State Transition Probabilities. The probability of transitioning from state $s_{i}^{t}$ to state $s_{j}^{t+1}$ given action $a_{k}^{t}$ is represented by $P_{i,j}^{k}=P(s_{t+1}=s_{j}^{t+1}|s_{t}=s_{i}^{t},a_{t}=a_{k}^{t}),$ where the complete state transition probability matrix given action $a_{k}^{t}$ would then be a matrix $\mathbf{P}^{k}$ where each entry $P_{i,j}^{k}$ represents the transition probability from state $s_{i}^{t}$ to state $s_{j}^{t+1}$.
Reward Structure. The immediate reward received after transitioning from state $s_{i}^{t}$ to $s_{j}^{t+1}$ upon taking action $a_{k}^{t}$ is given by $R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1})$. The expected reward for taking action $a_{k}^{t}$ in state $s_{i}^{t}$ is:
$R(s_{i}^{t},a_{k}^{t})=\sum_{j=1}^{n}P_{i,j}^{k}R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1}).$
Policy. A policy $\pi$ is a mapping from states to actions. An optimal policy $\pi^{*}$ maximizes the expected cumulative discounted reward from any starting state. The value of a state $s_{i}$ under policy $\pi$ is:
$V^{\pi}(s_{i})=R(s_{i},\pi(s_{i}))+\gamma\sum_{j=1}^{n}P_{i,j}^{\pi(s_{i})}V^{\pi}(s_{j}),$
where $\gamma$ is the discount factor for future rewards.
Optimal Policies under Bellman Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the following recursive Bellman optimality equation
$V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$
On top of this MDP-based policy derivation model, we are able to develop a DRL-based plug-and-play module for maintenance action recommendation using neural network-based function/policy approximation, by accommodating a number of state-of-the-art DRL algorithms. Algorithm 4 shows the details of leveraging different DRL algorithms, including DQN, SAC, PPO, for selecting optimal maintenance action $a$ through evaluating the performance of each policy. Details of hyperparameters setting and how to apply the algorithm to the real-world case are explained in Section V-C4.
Input: Exploration rate $\epsilon$ (for DQN, SAC), learning rate $\alpha$, discount factor $\gamma$ and policy update frequency (for PPO, SAC)
Output: Optimal action $a$ for maintenance
/* Initialization */
1 Initialize the policy network and optionally value/critic network(s) with the input hyperparameters;
2 Initialize target network(s) if applicable (e.g., for DQN, SAC);
3 Initialize replay buffer if applicable;
4 for each episode do
5       Initialize the starting state $\bm{s}$;
6       while $\bm{s}$ is not a terminal state do
             /* Action Selection */
7             if using DQN or SAC then
8                   Select action $a$ using an $\epsilon$-greedy policy or according to the current policy and exploration strategy;
9
10             else if using PPO then
11                   Select action $a$ according to the current policy;
12
13             end if
14            Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
             /* Policy and (or) Value Network Update */
15             if using DQN then
16                   Store transition $(\bm{s},a,r,\bm{s^{\prime}})$ in replay buffer;
17                   Sample mini-batch and update the network(s) using TD error;
18                   Periodically update target network;
19
20             else if using PPO then
21                   Accumulate data for policy update;
22                   Update policy using PPO;
23
24             else if using SAC then
25                   Store transition in replay buffer;
26                   Sample mini-batch and update policy and value networks using SAC update rules;
27                   Adjust temperature parameter for entropy maximization;
28
29             end if
            /* Exploration Rate Adjustment */
30             Adjust the exploration rate $\epsilon$ if applicable;
31             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
32
33       end while
      /* Policy Evaluation */
34       Periodically evaluate the performance of the current policy;
35
36 end for
return Optimal maintenance action $a$
Algorithm 4 A Deep Reinforcement Learning Algorithm for Maintenance Action Recommendation
V Case Study
In this section, we perform a case study on aircraft engine maintenance using NASA’s C-MAPASS dataset [28] to evaluate our proposed TranDRL framework.
V-A Dataset
The NASA’s C-MAPSS dataset [28] contains four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes. The training data comprises inputs from three operating conditions and 21 sensors’ measurements. For training, every engine data point serves as a sample paired with its RUL. In the experiment, we applied a piecewise linear degradation model [29] for RUL determinations. The last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
Figure 2: An Overview of the Temporal Fusion Transformer.
V-B Remaining Useful Life Prediction
The framework given in Algorithm 3 makes it suitable to incorporate the Transformer-based time-series event forecasting networks with other feature extraction and post-processing modules. In the case study, we adopt the Temporal Fusion Transformer (TFT) [30] in Algorithm 3 to utilizes the advantages of both recurrent neural networks and Transformers [31] into a unified encoder-decoder architecture. Compared to traditional time series forecasting models, e.g., Long Short-Term Memory networks (LSTMs) [32], TFT offers enhanced interpretability and the capability to deal with multiple prediction horizons. The structure of TFT is shown in the Figure 2 and the includes following components:
Encoder. The encoder transforms the raw time-series data into a high-dimensional space using convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
Temporal Attention. This specialized module let the model focus on different time-steps, facilitating the capture of dynamic temporal patterns and long-range dependencies.
Decoder. Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for forecasting subsequent steps.
Static Inputs. Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
Interpretable Components. To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
Evaluation Results. In our comparisons, Figures 3 to 6 demonstrate the TFT algorithm’s superior performance over NBeats [33] and DeepAR [34] in predicting the RUL for a individual machine, as evidenced by consistently higher RUL values. This suggests TFT’s enhanced capability in capturing complex, time-dependent patterns in sensor data, highlighting its robustness across various machine behaviors. In contrast, Figure 7 shows that while federated learning models generate reasonably good average RUL values across multiple machines, they may not capture the intricate details as effectively as single-machine models. However, Figures 8 to 11 show a significant bias when models trained on a single machine are applied to other machines, underscoring the challenge of model generalizability in diverse operational environments. This analysis illustrates the trade-off between the granularity and specificity of single-machine models and the broader applicability but potentially lower precision of federated learning approaches, highlighting the need for context-aware modeling strategies in predictive maintenance.
Figure 3: The RUL prediction result for machine FD001.
Figure 4: The RUL prediction result for machine FD002.
Figure 5: The RUL prediction result for machine FD003.
Figure 6: The RUL prediction result for machine FD004.
Figure 7: Comparison of RUL prediction results of four machines using the averaged model.
Figure 8: Comparison of RUL prediction results of four machines using the model trained from machine FD001.
Figure 9: Comparison of RUL prediction results of four machines using the model trained from machine FD002.
Figure 10: Comparison of RUL prediction results of four machines using the model trained from machine FD003.
Figure 11: Comparison of RUL prediction results of four machines using the model trained from machine FD004.
V-C Optimal Maintenance Actions Recommendation
In the RUL prediction stage, experiment results show that our proposed framework succeeds in predicting RULs using four datasets from NASA C-MAPSS dataset [28]. However, a gap still exists between the prediction results and the maintenance action decisions needed by engineers. To make action recommendation by weighing the cost and benefits, we leverage DRL algorithms as given in Algorithm 4, which helps to find the best trade-off.
V-C1 States and Actions
The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset. These PCA features provide a compact and important part of the machine’s current health condition. The action space contains three actions:
  • No Action ($a_{1}$), appropriate for healthy machine states, minimizes immediate costs but risks penalties in critical conditions;
  • Partial Maintenance ($a_{2}$), suitable for moderate unhealthy states, involves preventive actions like lubrication to extend machine life at moderate costs;
  • Complete Overhaul ($a_{3}$), necessary for critical machine states, entails comprehensive repairs and replacements to prevent catastrophic failures but incurs high costs and downtime.
The DRL agent aims to navigate this state-action space to minimize long-term operational costs by selecting the optimal maintenance actions based on each state’s representation of machine health.
V-C2 Probabilities Model
In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset. Specifically, each entry in the state transition matrix $p(S_{t+1}|S_{t},a_{t})$ reflects the probability of transitioning from state $S_{t}$ to $S_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset. Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset. In this MDP framework, the transition matrix has a dimension $(10,3,10)$, representing transitions between $10$ discretized states given $3$ actions, all closely reflecting scenarios that commercial aero-engines might encounter.
V-C3 Reward Model
The objective of maintenance actions recommend algorithms is not only to maximize the RUL but also to minimize the operational costs and downtime. To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner. The reward function $R(S_{t},a_{t},S_{t+1})$ is formulated as:
$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})~{}$
$-\beta\times\text{Cost}(a_{t})-\gamma\times\text{DownTime}(a_{t}),$ (6)
where
  • $\text{RUL}_{\text{gain}}(S_{t},a_{t})$ represents the gain in RUL as a consequence of taking action $a_{t}$ at state $S_{t}$. This gain is computed using the state-transition probabilities $p(S_{t+1}|S_{t},a_{t})$ derived from the C-MAPSS dataset. A higher RUL gain is desirable as it suggests an extension in the machine’s normal working time.
  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on experiences.
  • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$. Downtime is critical in operational settings, which impacts productivity or service levels.
  • $\alpha,\beta,\text{and}~{}\gamma$ are hyperparameters that weigh the importance of RUL, cost, and downtime, respectively. By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL, minimizing costs, or reducing downtime.
V-C4 Hyperparamters
For each DRL algorithm embedded in Algorithm 4, we carefully choose its corresponding hyperparameters based on empirical experiments. Details for each DRL algorithm are given as follows.
Deep Q-network (DQN). DQN reduces the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function [35]. It also introduces experience replay to improve learning stability and efficiency. The learning rate is set to $10^{-4}$, the discount factor is set to $0.99$, and the replay buffer size is set to $10^{6}$ experiences. We use a batch size of $512$ and update the target network every $50$ steps. The exploration-exploitation trade-off is controlled by an $\epsilon$ value that started at $1.0$ and decayed over time to a final value of $0.01$. Additionally, we use a neural network with two hidden layers of $256$ units each, and we utilize the Adam optimizer with a learning rate of $10^{-4}$.
Proximal Policy Optimization (PPO). PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy [36]. We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$. A clip range of $0.2$ is used to limit the updates to the policy, and we apply a value function coefficient of $0.5$ to balance the importance of the policy and the value function. Additionally, we set the entropy coefficient to $0.01$ to encourage exploration and a batch size as $64$. Then, we optimize the policy every 10 epochs and use the Adam optimizer with a learning rate of $0.0001$. Finally, GAE parameter is set to $0.95$ to estimate the advantage function.
Soft Actor-Critic (SAC). SAC is particularly effective for problems with high-dimensional action spaces [37]. SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient. The learning rate for the actor and critic networks is set to $0.0003$, and we use a discount factor of $0.99$. Then, we apply a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function. The target entropy is set to $-2.0$, and we use a replay buffer of $10^{6}$ experiences with a batch size of $256$. We also apply a soft update of the target networks with a rate of $0.005$, and we initialize the log standard deviation of the policy with a value of $-2$. Finally, we train the algorithm for $10^{5}$ time steps and use the Adam optimizer with a learning rate of $0.0001$.
V-D Results and Discussion
Figure 12: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
Figure 13: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
We compare the performance of DQN, PPO, and SAC on our designed task in Figures 1213. The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes. The experimental results show that PPO outperforms both DQN and SAC, achieving higher total rewards on the task. The average total rewards obtained by PPO are $55,000$, while DQN and SAC achieve average total rewards of $50,000$ and $45,000$, respectively. However, DQN converges faster than PPO and SAC.
In general, PPO and SAC are considered to be more sample-efficient than DQN, which means that they can achieve comparable performance with fewer samples. This is because PPO and SAC use more efficient updates that make better use of the available data. In contrast, DQN can be more sensitive to the choice of hyperparameters and the amount of data used during training, which can affect its convergence speed. However, the performance of these algorithms can vary depending on the specific task and the quality of the implementation. In some cases, DQN may converge faster than PPO or SAC, especially if the action space is discrete and the observations are low-dimensional. On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster.
Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve the optimal performance.
Figure 14: Differences between RL and RLHF.
V-E Reinforcement Learning with Human Feedback
Incorporating human expertise into the reinforcement learning process helps to resolve some of the challenges encountered with traditional RL methods. Especially, in scenarios where real-world rewards are hard to quantify or can be misleading, human feedback provides a way to guide the learning process more effectively. The comparison between RL and RLHF is presented in Fig. 14.
As machines and their associated systems become increasingly complex, the conditions that need to be considered maybe take a long time for Algorithm 5. Human experts, with their working experiences and domain knowledge to find complex patterns, can provide valuable insights that may not be readily available through raw data. By integrating human feedback into the RL process, the framework can:
  • Augment the learning process with domain-specific knowledge.
  • Address the cold start problem where initial exploration might be risky or expensive.
  • Improve learning efficiency by reducing the exploration of sub-optimal actions.
Thus, we further incorporate the following information into our proposed framework in Algorithm 5:
  1. 1.
    Human Opinions: Collect human predictions of the desired behavior. Experienced engineers can provide opinions on maintenance decisions based on various machine states from the C-MAPSS dataset.
  2. 2.
    Feedback on Agent’s Actions: After the agent makes a decision, a human can provide feedback on the chosen action, guiding the agent on the appropriateness of its actions.
Then, we introduce a term in the reward function that captures human feedback
$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if human feedback is positive},\\ r_{\text{negative}},&\text{if human feedback is negative},\\ 0,&\text{otherwise},\end{cases}$ (7)
based on which, the modified reward function for DRL is
$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})-\beta\times\text{Cost}(a_{t})-$
$\gamma\times\text{DownTime}(a_{t})+\delta\times R_{\text{human}}(a_{t}),$
where $\delta$ is a hyperparameter weighing the importance of human feedback.
Input: Observation $\bm{s}$, Exploration rate $\epsilon$, Learning rate $\alpha$, Discount factor $\gamma$, Human feedback 00.
Output: Optimal action $a$ for maintenance
1 Initialize the neural network weights;
2 Initialize a mechanism to incorporate human feedback (e.g., reward shaping, action recommendation);
3 for each episode do
       /* Incorporate Human Demonstrations */
4       If available, utilize human demonstrations to set initial policy;
5       Initialize the starting state $\bm{s}$;
6       while $s$ is not a terminal state do
             /* Action Selection with Human Feedback */
7             Select action $a$ using an $\epsilon$-greedy policy or based on human recommendations;
8             Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
             /* Adjust Reward Based on Human Feedback */
9             Incorporate human feedback into the reward:
$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if HF is positive},\\ r_{\text{negative}},&\text{if HF is negative},\\ 0,&\text{otherwise}.\end{cases}$
Adjust reward $r$ based on $R_{\text{human}}(a_{t})$;
             /* Network Update */
10             Calculate TD error $\delta=r+\gamma\max_{a^{\prime}}Q(\bm{s^{\prime}},a^{\prime})-Q(\bm{s},a)$;
11             Update neural network weights using $\delta$ and learning rate $\alpha$;
             /* Adjustment of Exploration and State */
12             Adjust the exploration rate $\epsilon$;
13             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
14
15       end while
      /* Incorporate Feedback into Learning Process */
16       Periodically adjust the learning process based on accumulated human feedback;
17
18 end for
return Optimal maintenance action $a$.
Algorithm 5 The RLHF Algorithm for Maintenance Actions Selection
V-F Results for RLHF
Figure 15: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
Figure 16: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
Figure 15 and Figure 16 reflect the changes after adding on the human feedback in the reward function. The results show that the speed of convergence does not improve too much, but the process becomes more stable and variances minimized. This is affected by the weight of human feedback in the reward function as well as the complexity of the task. Since our task converges very fast even without human feedback, there is not much room for improvement. In the future, we will conduct experiments on more complex real-world scenarios to confirm the effectiveness of the RLHF algorithm.
VI Conclusions
In this paper, we present a comprehensive prescriptive maintenance framework that integrates real-time RUL estimations with DRL algorithms for dynamic action recommendations. The framework starts by employing the Transformer network to estimate the RUL distribution of engines, offering a robust and adaptive data-driven approach to understanding machine conditions in real-time. As machine operates, our model continues to dynamically updating RUL estimates, providing an ever-evolving status of the equipment’s condition.
We rigorously compare the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates. Our findings indicate that the Transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process. For DRL-based maintenance action recommendation, we demonstrate the efficiency of SAC, DQN, and PPO. Each of these algorithms has its strengths and weaknesses, but they all are adapted to working seamlessly with the dynamically updated RUL estimates. Future work may involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machines.
In conclusion, our work signifies a step forward in the request for more intelligent, adaptive, and cost-effective prescriptive maintenance systems. By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a solution that dynamically adapts to the operational realities, thereby enhancing both economic and operational efficiencies.